OpenStudy (anonymous):

The equation of a curve is y = √ (5x + 4). (i) Calculate the gradient of the curve at the point where x = 1. [3] (ii) A point with coordinates (x, y) moves along the curve in such a way that the rate of increase of x has the constant value 0.03 units per second. Find the rate of increase of y at the instant when x = 1. [2] (iii) Find the area enclosed by the curve, the x-axis, the y-axis and the line x = 1. [5]

OpenStudy (gorv):

gradient mean \[\frac{ dy }{ dx}\]

OpenStudy (gorv):

first calculate the dy/dx

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 1 }{ 2 }\left( 5x+4 \right)^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

got that far

OpenStudy (gorv):

loll you forgot to differentiate 5x+4

OpenStudy (gorv):

after differentiating the root we also differentiate what is inside it

OpenStudy (gorv):

okk...got it ??..this will be our gradient

OpenStudy (gorv):

now rate of chnage of x = dx/dt

OpenStudy (gorv):

differentiate y with respect with t

OpenStudy (dumbcow):

use this \[\frac{dy}{dt} = \frac{dy}{dx}*\frac{dx}{dt}\]